what is the area of triangle?

Grade:7

13 Answers

Rahul Jiji George
9 years ago
Area of a triangle= ½ height * breadth
Ashwin Satheesh
9 years ago
AREA OF A RIGHT ANGLED TRIANGLE OR IN A TRIANGLE WHEN H IS GIVEN “THE AREA = ½ bh” . ALSO AREA OF A TRIANGLE WHEN ALL THE “sides” ARE GIVEN = SQUARE ROOT OF s×  (s-a)×(s-b)×(s-c)       HERE s = a+b+c/2    a,b,c are the sides of triangle.
Yugabrat Gogoi
9 years ago
area of triangle= ½ x base x hight. there is also a formula of heron, where area =under root s(s-a)(s-b)(s-c) where s= (a+b+c)/2 and where a, b, c are the sides of a triangle
Nihal Patel
9 years ago
Trinagle’s Area=1/2*bh
Triangle’s Area by Heron’s Formula= This formula works when three sides are given, which are a,b and c.First find the s.
=> s= a+b+c?/2
Then find,
s-a, s-b and s-c
Then, imply it in the main formula
s(s-a)(s-b)(s-c),
and the root of the obtained answer.
Thank You.
 
Ritika Das
6 years ago
Generally area of a triangle is equal to half of its product of its base length and height length. 
If the triangle is not a right angled triangle then we need to drop a perpendicular and calculate the resulting area by finding it. 
We also have Heron’s formula where if we know the perimeter and the individual side lengths, we can find the area by whole root over s*(s – a)*(s – b)*(s – c).
pranav kumar
6 years ago
when height and base are given then it is ½ into base into height
but when all the three sides are giiven then 
let one side be ‘a’ second be ‘b’ and third will be’c’
calculate the semi perimeter
semi perimeter is written as s
s=a+b+c
   ----------- 
      2
a+b+c whole divided by 2
then area=under root s(s-a)(s-b)(s-c)
thanks
please approve
 
 
swarup kumar majee
6 years ago
area of triangle is ½ base*height or 1/2*b*h
so,the main formula for area of triangle is ½ *base*height
Harshit Panwar
6 years ago
Area of triangle= 1/2(base×height)
Hope it helps you.
Priyansh
6 years ago
 Area of the triangle is given by mainly 4 ways 
 
1) base x height/2 ( Elementary Level)
2) ab x bc x Sin C (Trigonometry)
3) Heron's Formula ( Secondary Level)
4) 1/2 [x1 (y2-y3) +x2 (y3-y1) +x3 (y1-y2)] {Coordinate Geometry}
SR Roy
6 years ago
 
Trinagle’s Area=1/2*bh
Triangle’s Area by Heron’s Formula= This formula works when three sides are given, which are a,b and c.First find the s.
=> s= a+b+c?/2
Then find,
s-a, s-b and s-c
Then, imply it in the main formula
s(s-a)(s-b)(s-c),
and the root of the obtained answer.
 
 
Siddharth Singh
6 years ago
Area of triangle is =1/2 length*breadth........By herons formula ......={s*(s-a)*(s-b)*(s-c)}......where s =a+b+c/2
Om
5 years ago
Dear student
 
Area of triangle = ½ *base*height     (or)
You can even use herons formula to find the area of the triangle
 
I hope this answer helps you
 
With regards
A fellow student
Varsha Shivanand K
5 years ago
To find area of a triangle, we need a formula
Area of a triangle =1/2 * base * height
By using a formulae it is very easy to solve the problem.
Hope my answer helps
With regards,
Varsha Shivanand K.